# Computational Sentence-level Metrics Predicting Human Sentence Comprehension

### v1

@@ 4.1 Overall performance
We further compared the performance of the GAMM fittings with different metrics. Table 2 presents the results on $\Delta$AIC for comparing these GAMM fittings. Lower $\Delta$AIC value indicates a better GAMM fitting. It also suggests that the computational metric performs better other. The basis for comparison is the consistent data point numbers (n = 34571) and identical elements in each GAMM fitting. As illustrated in Table 2, the result reveals that sentence surprisal, as calculated using three distinct methodologies — namely, CR, NLL, and NSP — proves to be effective in predicting reading speed. Among the evaluated metrics of sentence surprisal, the performance of m-BERT, when applying the chain rule, stands out as the most effective. Conversely, sentence surprisal calculations using mGPT with NLL were found to be ineffective. The computation of sentence relevance, whether through m-BERT or mGPT, demonstrates viability in prediction accuracy. Notably, the integration of both sentence surprisal and sentence relevance into the GAMM significantly enhances predictive performance, surpassing the results achieved when these metrics are applied independently.
@@ 4.2 Performance in individual languages
The results from fitting the GAMM offer valuable insights into the importance of the variables considered. Using the $\Delta$AIC, we can determine the relative performance of different GAMM configurations. We identified the optimal metrics for our study: sentence surprisal calculated using m-BERT with the chain rule (CR), and sentence relevance determined by m-BERT. This analysis enables us to select optimal GAMM structures for evaluating how computational metrics predict sentence reading speed across 13 languages. The optimal GAMM for each language is specified as follows:reading_speed $\sim$ s (mean_word_length), s(mean_log_wordfreq) + s(sentence  surprisal) + s(sentence  relevance) + s (participant, bs=“re”), data= language_data. To calculate the $\Delta$AIC for “sentence relevance”, we excluded $s(sentence\quad surprisal)$ from the model, comparing the AIC of the complete model with that of the model absent “sentence surprisal”. Conversely, to determine the $\Delta$AIC for sentence surprisal,$s(sentence\quad relevance)$ is removed, and the AIC of the full model is subtracted from the model without “sentence relevance”. The significance of either “sentence surprisal” or “sentence relevance” is assessed using the p-value ( at p-value of 0.01) and the shape of the curve. These metrics were selected to represent, respectively, the concepts of sentence surprisal and sentence relevance in our examination of each language. The results are illustrated in Fig 4 in Appendix C. It appears that sentence surprisal lacks significance in English, Korean, and Russian, while sentence relevance shows no significant impact in English, Russian, Spanish, and Turkish.
We employed T-test to assess the statistical significance of the $\Delta$AIC values for both sentence surprisal and sentence relevance. The analyses revealed no significant differences, with the p-value substantially exceeding 0.05. This outcome suggests that sentence surprisal and sentence relevance may contribute equivalently to the prediction of reading speeds. Consequently, it can be inferred that both metrics are effective in facilitating the overall comprehension of sentences. Despite this, Table 2 shows that the overall predictive power of sentence surprisal with m-BERT and CR is stronger than sentence relevance. Moreover, sentence surprisal and sentence relevance are totally distinct metrics, and their overall correlation is -0.054 (p-value being zero), their correlation in each language is also remarkably small, seen in Appendix D.
@@ 5 Discussion
Word surprisal has been found to predict language comprehension across various datasets (Ryskin and Nieuwland, 2023). Sentence prediction has also been found in impacting language comprehension(Goldstein et al., 2022; Kriegeskorte et al., 2008). However, sentence-level computational metrics have not been proposed previously. Our study shows that the surprisal of a sentence can predict and clarify the difficulties readers encounter in comprehending a sentence as a whole. If a sentence is highly predictable given its left context, readers may feel easier to comprehend or process, leading to higher/faster reading speed. On the other hand, if a sentence is more unpredictable (i.e., lower sentence surprisal), readers may be more likely to experience processing difficulty. This can lead to slower/lower reading speed and potentially disrupt the flow of reading. The ability to predict the next sentence may be a fundamental aspect of language comprehension, as it allows users to anticipate upcoming information, and to integrate it smoothly with the preceding information. The impact of sentence surprisal on reading aligns closely with the influence of word surprisal on the reading process(Goodkind and Bicknell, 2018; Gotlieb Wilcox et al., 2023). Prediction occurs not only at the level of individual words, but also at the level of entire sentences in discourse. Prediction could be a key mechanism underlying human language comprehension, which involves using the context and prior knowledge to make predictions about the likely content of the upcoming input, and adjusting those predictions as new information is encountered.
@@
Numerous studies have also identified the influence of context on sentence processing (Cohen and Servan-Schreiber, 1992; Grisoni et al., 2017). However, we are the first to propose to compute sentence-level semantic relevance. The sentence relevance computed by attention-aware method can also effectively help interpret and predict the processing difficulties readers face when comprehending sentences in their entirety. When a sentence is more semantically relevant to the context (in discourse) in which it appears, it is more likely to be processed quickly and accurately, and readers are more likely to understand its meaning without having to invest a lot of cognitive effort (i.e. higher reading speed). Conversely, when a sentence is irrelevant with the context (i.e., lower sentence relevance), it may slow down reading speed and require more cognitive effort. The impact of sentence relevance on reading closely resembles the effect of word semantic relevance on the reading (Sun et al., 2023b; Sun et al., 2023a). There are several reasons why relevant sentences tend to be processed more quickly. According to the research of discourse structure, relevant sentences can easily create discourse coherence. One key factor is the activation of mental representations related to the topic or theme of the text (Traxler, 2011). When a sentence is relevant to the context, it could activate mental representations through memorizing the context, making it easier for readers to process the sentence. This activation can lead to faster reading speed, as readers are able to effectively activate the memory of the context and quickly integrate it into their mental representation of the text. In contrast, irrelevant sentences may require more cognitive effort to process, as they could not activate the memory from the context and do not fit smoothly into the mental representation of the discourse that readers are constructing. This processing difficulty can slow down reading speed.
@@
The expectation and memory could be mutually interacted (Ryskin and Nieuwland, 2023) in language comprehension/processing. Sentence surprisal measures the processing difficulty of a sentence by assessing its unexpectedness relative to the preceding context. On the other hand, sentence relevance represents memory by gauging how well a sentence fits into the context, with contextually relevant sentences being easier to process, resulting in faster reading speed. By combining them, it can encompass both processing difficulty, and could better simulate human language processing.
@@ 6 Conclusion
This study presented sentence-level metrics for predicting human comprehension of sentences as a whole. The results show that both sentence surprisal and sentence relevance were highly capable of predicting human sentence prediction. The methods of computing sentence surprisal worked well. The attention-aware method allowed for computing contextual information for sentence-level semantic relevance. All of these methods also exhibited strong generalization capabilities across languages. The findings showed that these sentence-level metrics are informative features for modeling human sentence comprehension across languages. Our work highlighted the potential of combining computational models with cognitive models to better explain and predict human language comprehension/processing, and further to develop more effective NLP and AGI systems.
@@ 7 Limitations
The study introduces an innovative approach to predicting sentence comprehension using attention-aware computational metrics. However, it comes with several limitations. Primarily, the methodology depends on specific multilingual LLM (m-BERT and mGPT). The variation in proficiency across different languages can result in disparate abilities to process these languages. In essence, sentence surprisal or sentence relevance might not be estimated with equal precision across all languages using multilingual LLMs. This limitation could affect the applicability of our metrics across varied linguistic structures and cultural contexts, leading to varying performance of our metrics in different languages. Moreover, the attention-aware method, while effectively capturing some aspects of contextual information, may oversimplify the complex dynamics of human cognitive mechanisms on language comprehension and processing. Addressing these limitations could pave the way for more comprehensive models that better mirror the intricacies of human language comprehension and processing.
@@
A. Computing sentence surprisal
@@
The following provides a detailed account of how to apply chain rule (CR) to compute sentence surprisal . First, we tokenized sentence into various tokens: $S=[t_{1},t_{2},\ldots,t_{n}]$. Probability of a token $t_{i}$ in the sentence S given its left context C in a text and preceding tokens can be represented as: $P(t_{i}|C,t_{1},t_{2},\ldots,t_{i-1})$. (S = sentence, t = token, C = context). Based on this, we can calculate the probability of each token prior to its left context. After obtaining the probability of each word, we computed the joint probability of the entire sentence in equation (2): S given its context C, $P(S|C)=P(t_{1}|C)\cdot P(t_{2}|C,t_{1})\cdot P(t_{3}|C,t_{1},t_{2})\cdot\ldots\cdot P(t_{n}|C,t_{1},t_{2},\ldots,t_{n-1})\quad(2)$, further getting sentence surprisal, $-log(P(S|C))$. The equation is an application of the chain rule of probability to sequences (like sentences). It captures the idea that the probability of a sentence given a context can be decomposed into the product of conditional probabilities of its individual tokens. The following provides the details why it works. When you have a sequence of events or tokens, the chain rule can be extended. For a sequence of three events $t_{1}$, $t_{2}$, and $t_{3}$: $P(t_{1},t_{2},t_{3})=P(t_{1})\times P(t_{2}|t_{1})\times P(t_{3}|t_{1},t_{2})$. In the case of natural language, sentences often have a context. The context can be prior knowledge, a preceding sentence, or any other relevant information. Given a context (C), the probability of a sequence changes (S): $P(S|C)=P(t_{1}|C)\times P(t_{2}|C,t_{1})\times P(t_{3}|C,t_{1},t_{2})\times\ldots\times P(t_{n}|C,t_{1},t_{2},...,t_{n-1})$. The method of computing sentence probability is consistent with the principles of probability theory and the way sequence modeling is approached in the context of NLP.
@@
In addition to adopting the chain rule, we could employ the “next sentence prediction (NSP)” mechanism in BERT to compute sentence surprisal. To compute the probability that one sentence follows another using BERT’s NSP, begin by preparing the input. Combine the two sentences, placing a special [CLS] token at the start and a [SEP] token in between them. After tokenizing this combined sequence with BERT’s tokenizer, pass it through the BERT model. BERT uses the representation of the [CLS] token, which encapsulates information about the entire input, to predict the relationship between the two sentences. The model outputs two probabilities: “IsNext” and “NotNext”. The “IsNext” probability indicates the likelihood that the second sentence logically follows the first. By examining this probability, one can gauge how likely the model perceives the given sentence order to be. The above description can be summarized as the following two equations. Probability of sentence B being the next sentence given the left context A is formalized as (4):
@@
where [CLS] denotes the representation of the first sentence obtained from the final layer of the BERT model, and the softmax function computes the probability distribution over all possible next sentences. $w_{2}$ and $b_{2}$ are learnable weight vector and bias term, respectively. And then we can obtain the sentence surprisal. However, NSP in BERT is a binary classifier, that is, the model determines if a sentence logically follows a given sentence or not. This is a simplification of the rich structure and semantics in natural language. The use of NSP may raise the question about the interpretability of BERT’s estimated probabilities for subsequent sentences, especially when viewed from a cognitive modeling standpoint.
B. Attention-aware approach and its memory capability

### v2

@@ 4.1 Overall performance
We further compared the performance of the GAMM fittings with different metrics. Table 2 presents the results on $\Delta$AIC for comparing these GAMM fittings. Lower $\Delta$AIC value indicates a better GAMM fitting. It also suggests that the computational metric performs better than others. The basis for comparison is the consistent data point numbers (n = 34571) and identical elements in each GAMM fitting. As illustrated in Table 2, the result reveals that sentence surprisal, as calculated using three distinct methodologies — namely, CR, NLL, and NSP — proves to be effective in predicting reading speed. Among the evaluated metrics of sentence surprisal, the performance of m-BERT, when applying the chain rule, stands out as the most effective. Conversely, sentence surprisal calculations using mGPT with NLL were found to be ineffective in predicting reading speed. We therefore skiped the result of mGPT with NLL. The computation of sentence relevance, whether through m-BERT or mGPT, demonstrates viability in prediction accuracy. Notably, the integration of both sentence surprisal and sentence relevance into the GAMM significantly enhances predictive performance, surpassing the results achieved when these metrics are applied independently.
@@ 4.2 Performance in individual languages
The results from fitting the GAMM offer valuable insights into the importance of the variables considered. This section focuses on the performance of these metrics in each language (among 13 languages) using the similar GAMM analysis. The optimal GAMM for each language is specified as follows:reading_speed $\sim$ s (mean_word_length), s(mean_log_wordfreq) + s(sentence  surprisal) + s(sentence  relevance) + s (participant, bs=“re”), data= language_data. To calculate the $\Delta$AIC for “sentence relevance”, we excluded $s(sentence\quad surprisal)$ from the model, comparing the AIC of the complete model with that of the model absent “sentence surprisal”. Conversely, to determine the $\Delta$AIC for sentence surprisal,$s(sentence\quad relevance)$ is removed, and the AIC of the full model is subtracted from the model without “sentence relevance”. The significance of either “sentence surprisal” or “sentence relevance” is assessed using the p-value ( at p-value of 0.01) and the shape of the curve. These metrics were selected to represent, respectively, the concepts of sentence surprisal and sentence relevance in our examination of each language. The results are illustrated in Fig. 5 in Appendix F. It turns out that sentence surprisal and sentence relevance shows significance in the majority of languages. Moreover, sentence surprisal and sentence relevance are totally distinct metrics, and their overall correlation is -0.054, their correlation in each language is also remarkably small. This indicates that sentence surprisal and sentence relevance are distinct metrics, and more are seen in Appendix G.
@@ 5 Discussion
Word surprisal has been found to predict language comprehension across various datasets (Ryskin and Nieuwland, 2023). Sentence prediction has also been found in impacting language comprehension(Goldstein et al., 2022; Kriegeskorte et al., 2008). However, sentence-level computational metrics have not been proposed previously. Our study found that the surprisal of a sentence can predict and clarify the difficulties readers encounter in comprehending a sentence as a whole. Specifically, if a sentence is highly predictable given its left context, readers may feel easier to comprehend or process, leading to higher/faster reading speed. On the other hand, if a sentence is more unpredictable (i.e., lower sentence surprisal), readers may be more likely to experience processing difficulty. This can lead to slower/lower reading speed and potentially disrupt the flow of reading. The ability to predict the next sentence may be a fundamental aspect of language comprehension, as it allows users to anticipate upcoming information, and to integrate it smoothly with the preceding information. The impact of sentence surprisal on reading aligns closely with the influence of word surprisal on the reading process(Goodkind and Bicknell, 2018; Gotlieb Wilcox et al., 2023). Prediction occurs not only at the level of individual words, but also at the level of entire sentences in discourse. Prediction could be a key mechanism underlying human language comprehension, which makes predictions about the likely content of the upcoming input.
@@
Numerous studies have also identified the influence of context on sentence processing (Cohen and Servan-Schreiber, 1992; Grisoni et al., 2017). However, we are the first to propose to compute sentence-level semantic relevance. The sentence relevance computed by the “attention-aware” method can effectively help interpret and predict the processing difficulties readers face when comprehending sentences in their entirety. When a sentence is more semantically relevant to the context (in discourse) in which it appears, it is more likely to be processed quickly and accurately, and readers are more likely to understand its meaning without having to invest a lot of cognitive effort (i.e. higher reading speed). Conversely, when a sentence is irrelevant with the context (i.e., lower sentence relevance), it may slow down reading speed and require more cognitive effort. The impact of sentence relevance on reading closely resembles the effect of word semantic relevance on the reading (Sun et al., 2023b; Sun et al., 2023a). There are several possible reasons why relevant sentences tend to be processed more quickly. According to the research of discourse structure, relevant sentences can easily create discourse coherence. One key factor is the activation of mental representations related to the topic or theme of the text (Traxler, 2011). When a sentence is relevant to the context, it could activate mental representations through memorizing the context, making it easier for readers to process the sentence. This activation can lead to faster reading speed, as readers are able to effectively activate the memory of the context and quickly integrate it into their mental representation of the text. In contrast, irrelevant sentences may require more cognitive effort to process, as they could not activate the memory from the context and do not fit smoothly into the mental representation of the discourse that readers are constructing. This processing difficulty can slow down reading speed.
@@
The expectation and memory could be mutually interacted (Ryskin and Nieuwland, 2023) in language comprehension/processing. By combining sentence surprisal and sentence relevance (shown in Table 2), it can encompass both processing difficulty for an entire sentence, and could better simulate human language processing.
@@ 6 Conclusion
This study presented two sentence-level metrics for predicting human comprehension of sentences as a whole. The methods of computing sentence surprisal worked well, and the attention-aware method allowed for computing contextual information for sentence-level semantic relevance. The results show that both sentence surprisal and sentence relevance were highly capable of predicting human sentence reading speed. All of these methods also exhibited strong generalization capabilities across languages. The findings showed that these sentence-level metrics are informative features for modeling entire sentence comprehension by humans across languages. Our work highlighted the potential of combining computational models with cognitive models to better explain and predict human language comprehension/processing, and further to develop more effective NLP and AGI systems.
@@ 7 Limitations
The study introduces innovative approaches to predicting entire sentence comprehension using computational sentence-level metrics. However, it comes with several limitations. Primarily, the methodology depends on specific multilingual LLM (m-BERT and mGPT). The variation in proficiency across different languages can result in disparate abilities to process these languages. In essence, sentence surprisal or sentence relevance might not be estimated with equal precision across all languages using multilingual LLMs. This limitation could affect the applicability of our metrics across varied linguistic structures and cultural contexts, leading to varying performance of our metrics in different languages. Moreover, the attention-aware method, while effectively capturing some aspects of contextual information, may oversimplify the complex dynamics of human cognitive mechanisms on language comprehension and processing. Addressing these limitations could pave the way for more comprehensive models that better mirror the intricacies of human language comprehension and processing.
@@
Appendices
A. Differences between reading speed and fixation duration
In reading research, reading speed and fixation duration for individual words represent distinct but related aspects of how we process language. The following specifies these differences.
Reading speed (or average reading speed, or reading rate) is typically measured in words per minute or second (e.g., wpm) and reflects the overall rate at which a person reads a passage of text or a sentence. Reading speed is usually employing text or individual sentence as the unit of interest. This metric gives a general indication of reading efficiency or fluency. It measures the reader’s ability to comprehend and process text or sentence over a given period. This can vary significantly depending on the reader’s skill, the text’s complexity, and the reading purpose (e.g., reading for pleasure vs. critical reading). Reading speed encompasses the entire reading process, including both the time spent on fixations (when the eyes stop on a word) and saccades (the rapid movements between fixations), as well as regressions (backward movements to re-read text) and pauses for comprehension. Reading speed has been extensively and intensively explored (Jackson and McClelland, 1979; Herman, 1985; Carver, 1990; Rayner et al., 2010).
Fixation duration for individual words (here total fixation duration) refers to the total length of time the eyes remain stationary during fixations on a word while reading. These durations are measured in milliseconds with an eye-tracker and can indicate the cognitive effort required to process the word. Fixation duration on individual words can reveal how word characteristics (such as length, frequency, and predictability) and contextual difficulty affect reading. Longer fixations tend to suggest that a word is harder to recognize or comprehend within its context. Fixation durations vary across words within a text and among different readers. They provide insight into the moment-by-moment cognitive processes involved in lexical reading, such as lexical access, parsing, and integration of information.
Reading speed offers a macro-level view of reading behavior, indicating overall efficiency or difficulty at the sentence or text level. In contrast, fixation duration provides a micro-level perspective, revealing the cognitive processing time for individual words. While both metrics can be affected by the text’s difficulty and the reader’s proficiency, reading speed could be also influenced by broader factors like braoder context and the reader’s strategy or purpose. Fixation duration is more directly related to immediate cognitive demands for word-specific characteristics.
B. Computing sentence surprisal
@@
The following provides a detailed account of how to apply chain rule (CR) to compute sentence surprisal . First, we tokenized sentence into various tokens: $S=[t_{1},t_{2},\ldots,t_{n}]$. Probability of a token $t_{i}$ in the sentence S given its left context C in a text and preceding tokens can be represented as: $P(t_{i}|C,t_{1},t_{2},\ldots,t_{i-1})$. (S = sentence, t = token, C = context). Based on this, we can calculate the probability of each token prior to its left context. After obtaining the probability of each word, we computed the joint probability of the entire sentence in equation (2): S given its context C, $P(S|C)=P(t_{1}|C)\cdot P(t_{2}|C,t_{1})\cdot P(t_{3}|C,t_{1},t_{2})\cdot\ldots\cdot P(t_{n}|C,t_{1},t_{2},\ldots,t_{n-1})\quad(2)$, further getting sentence surprisal, $-log(P(S|C))$. The equation is an application of the chain rule of probability to sequences (like sentences). It captures the idea that the probability of a sentence given a context can be decomposed into the product of conditional probabilities of its individual tokens. The following provides the details why it works. When you have a sequence of events or tokens, the chain rule can be extended. For a sequence of three events $t_{1}$, $t_{2}$, and $t_{3}$: $P(t_{1},t_{2},t_{3})=P(t_{1})\times P(t_{2}|t_{1})\times P(t_{3}|t_{1},t_{2})$. In the case of natural language, sentences often have a context. The context can be prior knowledge, a preceding sentence, or any other relevant information. Given a context (C), the probability of a sequence changes (S): $P(S|C)=P(t_{1}|C)\times P(t_{2}|C,t_{1})\times P(t_{3}|C,t_{1},t_{2})\times\ldots\times P(t_{n}|C,t_{1},t_{2},...,t_{n-1})$ (i.e., the multiplication of probabilities of each token given the left context within a sentence). The method of computing sentence probability is consistent with the principles of probability theory and the way sequence modeling is approached in the context of NLP.
@@
The following elaborates on the differences between applying Chain Rule (CR) and Negative Log Likelihood (NLL) within the contexts of BERT and GPT. We detail these distinctions specifically in the scenarios of BERT and GPT, respectively.
Within the context of autoregressive language models (LMs) like GPT, the concepts of the CR for probability and NLL indeed converge when computing sentence surprisal. The CR decomposes the probability of a sequence into the product of conditional probabilities. This method aligns with how autoregressive LMs generate text, predicting each token based on the previous ones. NLL is often used as a loss function in training these models, calculated by taking the negative logarithm of the probability of the observed data under the model. For sentence, NLL would be the sum of the negative logs of these conditional probabilities. When you take the logarithm of the product of conditional probabilities (as per the CR), it becomes a multiplication of these probabilities. The math equations for CR and NLL introduced at the start of this section are applicable to the GPT case.
However, in the context of BERT, we need to make slight changes because BERT is fundamentally different from autoregressive models in that it does not generate text sequentially. BERT’s architecture is designed to understand the context of a token by looking at the tokens that come before and after it—this is what makes it “bidirectional”. It is trained primarily through tasks like Masked Language Modeling (MLM), where it predicts the identity of tokens that have been artificially masked in the input text, based on the context provided by the non-masked tokens around them. The chain rule (CR), which breaks down the probability of a sequence into the product of conditional probabilities, naturally applies to models that generate or predict sequences in a linear, step-by-step manner (e.g., predicting the next word given the previous words in a sentence). While BERT does not compute sentence surprisal through the chain rule in the same way as autoregressive models, we can use BERT to estimate a form of surprisal or predictability for words in a sentence based on its bidirectional context. For instance, by masking a word in a sentence and using BERT to predict its likelihood, one can infer a measure of how predictable that word is given its context. BERT’s capacity to predict the probability of masked tokens in context allows for an indirect assessment of word-level surprisal. This approach leverages BERT’s contextual predictions to gauge how expected or surprising a word is within its sentence context. Assuming a sequence of tokens S = [$t_{1}$, $t_{2}$, …, $t_{n}$], and a context C, we conceptually mask and predict each token $t_{i}$ based on all other tokens as context, mimicking a sequential dependence that does not naturally exist in BERT. The “conditional probability” of each token $t_{i}$, given its context (both preceding and succeeding, C), is estimated as $P(t_{i}|C,t_{\text{rest}})$, where $t_{r}est$ represents the rest of the tokens serving as context which consists the target sentence and the previous stentences. Sentence surprisal would then be represented as Equation (4):
$\text{Surprisal}(S|C)=-\log\prod_{i=1}^{n}P(t_{i}|C,t_{\text{rest}})\qquad(5)$
We then focus on NLL in BERT. While NLL is a concept that can be applied broadly in machine learning to measure how well a model’s predictions match the actual data (and is used as a loss function in many contexts, including training BERT), its specific application and interpretation can vary. In the context of BERT, NLL is used to optimize the model’s ability to accurately predict the masked tokens based on their context, which involves both preceding and succeeding tokens in a sentence. In other words, BERT predicts the identity of masked tokens based on their surrounding (bidirectional) context, and NLL quantifies the model’s prediction accuracy. For each masked token prediction, compute the NLL as: $\text{NLL}(S|C)=-\sum\log(P(t_{i}|t_{1},t_{2},...,t_{i-1},t_{i+1},...,t_{n}))$ Here, $t_{i}$ is the masked token, and the context $C$ includes both preceding and succeeding tokens. Sum up the NLLs for all masked tokens in the sentence to get the total NLL, representing the sentence surprisal.
@@
In addition to adopting the chain rule and NLL, we could employ the “next sentence prediction (NSP)” mechanism in BERT to compute sentence surprisal. To compute the probability that one sentence follows another using BERT’s NSP, begin by preparing the input. Combine the two sentences, placing a special [CLS] token at the start and a [SEP] token in between them. After tokenizing this combined sequence with BERT’s tokenizer, pass it through the BERT model. BERT uses the representation of the [CLS] token, which encapsulates information about the entire input, to predict the relationship between the two sentences. The model outputs two probabilities: “IsNext” and “NotNext”. The “IsNext” probability indicates the likelihood that the second sentence logically follows the first. By examining this probability, one can gauge how likely the model perceives the given sentence order to be. The above description can be summarized as the following two equations. Probability of sentence B being the next sentence given the left context A is formalized as (6):
@@
where [CLS] denotes the representation of the first sentence obtained from the final layer of the BERT model, and the softmax function computes the probability distribution over all possible next sentences. $w_{2}$ and $b_{2}$ are learnable weight vector and bias term, respectively. And then we can obtain the sentence surprisal. However, NSP in BERT is a binary classifier, that is, the model determines if a sentence logically follows a given sentence or not. This is a simplification of the rich structure and semantics in natural language. The use of NSP may raise the question about the interpretability of BERT’s estimated probabilities for subsequent sentences, especially when viewed from a cognitive modeling standpoint. The reason for this is that we probably argue that as humans read, they probably take more than two sentences as the context window. Despite this, we still need to apply the statistical analysis to test whether such metrics are useful or not.
C. BERT approximates human reading behavior
Some researchers have underestimated BERT’s potential in calculating word probabilities, arguing that the use of BERT (masked LMs) is hard to justify within the realm of cognitive modeling. Their concern stems from the fact that a masked language model like m-BERT can consider context from both directions, including rightward context, which could not be typically how human comprehension works. However, BERT is better able to estimate word probability because it can be really line with human reading behavior. The following details this point.
Those familiar with eye-tracking in reading recognize the preview benefit (i.e., parafoveal-on-foveal effects, “PoF” for short). Parafoveal-on-foveal effects in reading refer to the influence of information from the words located in the parafoveal region (the area immediately surrounding the fixated word) on the processing of the currently fixated word (the foveal word) (Kennedy and Pynte, 2005; Kliegl et al., 2007; Schotter et al., 2012; Hohenstein and Kliegl, 2014). As the parafoveal-on-foveal effect takes place in the context of reading and visual processing, the visual perception of words not directly fixated upon (parafoveal words) influences the processing of the word currently fixated on (the foveal word).
Put it simply, when we read, our eyes do not move smoothly across the text but make quick, jerky movements known as saccades, interspersed with brief pauses called fixations. During these fixations, the eye directly focuses on a small area of text. This area, where visual acuity is highest, is known as the foveal region. Surrounding the foveal region is the parafoveal region, where vision is not as sharp but still capable of processing some information about the text, such as word length or initial letters. The PoF effect refers to how information from the parafoveal region can pre-activate or facilitate the processing of the word once it becomes the focus of the next fixation (moves into the foveal region), thereby affecting reading speed and comprehension. For example, if the parafoveal word is semantically or syntactically related to the foveal word, it might speed up the recognition and processing of the foveal word when the eye moves to it. In short, processing the target word actually is involved in incorporating the information on the next one or two ore more words (n+1, or n+2 word in literature) through previewing, as shown in Fig. 3.
Figure 3: The parafoveal-on-foveal effects in reading (from Sakurai (2023))
Relevant research shows that reading times are shorter for a target word when it matches the preview word, compared to when they are different. This suggests that the preview word is processed using parafoveal vision. BERT mirrors this process to some extent because it can incorporate information from subsequent words, approximating how human parafoveal vision previews words to the right of the fixation. In this sense, BERT aligns with human reading patterns to a greater degree. The surprisal values calculated by BERT, reflecting real-world reading dynamics, potentially provide more precise predictions of word reading times. The criticism that BERT generates only pseudo-surprisal overlooks its relevance to real-world human reading behaviors. On the other hand, surprisal values computed by GPT do not account for the previewing aspect of reading. Despite this, surprisal values from both GPT and BERT can, in theory, offer insights into reading behaviors. Nevertheless, statistical analysis is needed to determine which model’s surprisal predictions align more closely with observed reading data.
D. Attention-aware approach and its memory capability
This section provides a comprehensive guide on calculating sentence relevance using an attention-aware approach, executed in a two-step strategy. The sliding window includes four sentences (t, 2, 1, and n1), as shown in the Panel B of Fig. 1. “t” represents the target sentence, and “n1” for the next sentence. We aim to compute how the target sentence (t) is semantically related with the other three sentences (2, 1, and n1 form the context window).